OpenStudy (anonymous):

write the equation of a line that is perpendicular to the given line and that passes through the given point. -x+5y=14; (-5,-2)

OpenStudy (anonymous):

do you know the point-slope formula?

OpenStudy (anonymous):

yes I do. I'm up to y-(2)=5x-25 I don't know what to do next.

OpenStudy (amorfide):

get into the form y=mx+b then perpendicular would be the negative reciprocal of the slope m and then you use y-y1=m(x-x1) using x1 and y1 to be the coordinates given

OpenStudy (anonymous):

@amorfide I know this. I'm stuck at one part.

OpenStudy (anonymous):

the first step is to first find the slope of the given line. To do this, put it in slope intercept form: -x+5y=14 5y=x+14 y=1/5x+14/5 So the slope of this line is 1/5. If it is perpendicular to this, you know the slope will be the negative inverse, which would be -5 then plug what you know into the equation y-y1=m(x-x1) y-(-2)=-5(x+5) y+2= -5x-25 y=-5x-27

OpenStudy (anonymous):

where the heck does the 27 come from?

OpenStudy (anonymous):

you have to put it back into slope intercept form from y+2=-5x-25, you have to subtract 2 from each side to isolate y

OpenStudy (amorfide):

y-y1=m(x-x1) x1=-5 y1=-2 m=-5 y--2=-5(x--5) y+2=-5(x+5) y+2=-5x-25 since we expanded the brackets so just subtract 2 from both sides to get y=-5x-27 this is the step by step with an added step graze missed out you should understand it now hopefully